# Two-step estimation of latent trait models

### v2

@@ 3.5 Two-step estimation for other types of models
Two-step estimation could in principle be applied to any instance of the general latent variable model that was defined at the start of Section 2. To provide a focus for this paper, we concentrate on the family of latent trait models that was introduced later in Section 2. Before proceeding to empirical investigations of them, we comment briefly on other settings where the two-step method could be used. Some of them are obvious extensions or variants of the models considered here, while others could involve more consequential differences. Implementation and performance of two-step estimation in many of these other cases remain to be examined.
In the measurement model we made the assumption that $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})$. Relaxing this would take us to models with non-equivalence of measurement with respect to the observed covariates $\mathbf{X}_{i}$. Two-step estimation in that situation can introduce new issues, and deserves further investigation. For latent class models, it has been studied by Di Mari and Bakk (2018), Janssen et al. (2019) and Vermunt and Magidson (2021a).
@@
We made the common assumption that all items in $\mathbf{Y}_{i}$ are conditionally independent given the latent traits $\boldsymbol{\eta}_{i}$. This could be relaxed by allowing conditional associations between some items, requiring only corresponding modifications of the log likelihoods (9) and (10). Another obvious class of extensions would be obtained by considering models with mixtures of different types of measurement items. In this paper we focus on categorical items, while linear structural equation models (as in Rosseel and Loh 2022) take the items to be continuous. But some applications may involve both of these types, or others such as counts or survival times as measurement items. Similarly, the structural model could include both categorical and continuous latent variables, and categorical variables could be measured not only by categorical items (as in standard latent class analysis discussed by Bakk and Kuha 2018) but also by continuous ones (the case known as latent profile analysis). It seems plausible that the behaviour of the two-step method would not be fundamentally different in these different combinations of settings.
@@
Generalisations of the structural model could also include such possibilities as interactions or nonlinear terms of the latent variables, or non-recursive models. Two-step estimation in these cases has not been examined. It could behave differently there, if only because such models are inherently more complex and difficult howsoever they are estimated.
@@ 4.1 First set of simulations
We have used simulation studies to examine the properties of two-step estimates for the types of models considered in this paper. In this section we describe them for one model setting in detail. In Section 4.2 we then describe further simulations in several other settings.
The properties that we examine include the performance of the point estimates and accuracy of the estimated standard errors. The main comparison of interest is with one-step estimates of the same models. Since they are full maximum likelihood estimates, they serve as a benchmark whose theoretical properties are well understood. In large samples one-step estimates should be unbiased and fully efficient, and it is then interesting to see if two-step estimates do roughly as well. In smaller samples, either approach could have an advantage. In this section all the models are correctly specified, but in the next we also include simulations with a misspecified measurement model in order to examine the robustness of the estimates in such cases. Another point of comparison is computational demands, where two-step estimation should have the advantage. The simulation results also include (naive) three-step estimates, but they are not discussed here. We will comment on the results for them separately in Section 4.3.
@@
where $\eta_{1i}$ and $\eta_{2i}$ are latent variables and $\epsilon_{i}$ is the model residual. Here $\eta_{1i}\sim N(0,1)$ and $\epsilon_{i}\sim N(0,\sigma^{2}_{\epsilon})$, $\beta_{0}=0$, and $(\beta_{1},\sigma^{2}_{\epsilon})$ are such that marginally $\eta_{2i}\sim N(0,1)$ in all simulation settings. The value of $\beta_{1}$ is set so that the $R^{2}$ statistic for the response $\eta_{i2}$ is $R^{2}=0,0.2$, or $0.4$. We refer to this as simulation case A. Path diagram for this model is shown in Figure A.1 in the supplementary appendix.
The latent variable $\eta_{i1}$ is measured by $p$ binary items $Y_{i1j}$, coded with values 0 and 1, and $\eta_{i2}$ by a different set of $p$ binary items similarly, with $p=4$ or $p=8$. All items are conditionally independent of each other given the latent variables. The measurement model is of the form $\text{logit}[P(Y_{ikj}=1|\eta_{ik})]=\tau_{kj}+\lambda_{kj}\eta_{ik}$ for $k=1,2$ and $j=1,\dots,p$. The measurement parameters $(\tau_{kj},\lambda_{kj})$ are estimated as distinct parameters for different $k,j$, but their true values $\tau_{kj}=\tau$ and $\lambda_{kj}=\lambda$ in the data-generating model are set to be equal for all items in a given simulation. The loading parameter $\lambda$ is set with reference to the linear model for a notional continuous latent variable $Y_{ikj}^{*}$ which implies the logistic model for $Y_{ikj}$. Here $R^{2}$ for $Y_{ikj}^{*}$ given $\eta_{ik}$ is $R^{2}_{Y}=\lambda^{2}\mbox{var}(\eta_{ik})/(\lambda^{2}\mbox{var}(\eta_{ik})+\pi^{2}/3)$ where $\mbox{var}(\eta_{ik})=1$, and $\lambda$ is set so that $R^{2}_{Y}=0.4$ or 0.6. The values of $\tau$ are then set so that the marginal probability $\pi_{Y}=P(Y_{ikj}=1)$ is 0.5 or (approximately) 0.8. When the model is estimated, we constrain $\tau_{k1}=0$ and $\lambda_{k1}=1$ for the first item $Y_{ik1}$ of each latent variable, leaving $\tau_{kj}$ and $\lambda_{kj}$ for the other $p-1$ items estimable. This affects the implied scales of the latent variables, so that the value of $\beta_{1}$ that the estimators should actually be estimating will depend also on the value of $\lambda$ in the data-generating model. These true values of $\beta_{1}$ are shown in the results tables.
In each setting we generate 1000 simulated datasets with $n=200$ or $n=1000$ independent observations $i$. All combinations of $n$, $R^{2}$, $p$, $\pi_{Y}$ and $R^{2}_{Y}$ are considered, resulting in 48 settings. The two-step, one-step and naive three-step estimates are calculated. Two-step estimation was implemented as described in Section 3.3. The same R functions we wrote for it (combined with Mplus) also produced the one-step and three-step estimates.
@@ 4.2 Further simulations
  • Ordinal measurement items (case D). Structural model is the same as in case A, but each measurement item $Y_{ikj}$ has 4 rather than 2 categories. Measurement models are again derived from a linear model for a latent continuous $Y_{ijk}^{*}$, here resulting in an ordinal logistic measurement model for each $Y_{ijk}$. Its loading parameter $\lambda$ is set as in case A. The measurement intercepts are such that the marginal probabilities of the four categories of each $Y_{ikj}$ are (0.25,0.25,0.25,0.25) when $R^{2}_{Y}=0.6$ and about (0.20,0.30,0.30,0.20) when $R^{2}_{Y}=0.4$ (these correspond to the setting $\pi_{Y}=0.5$ in case A).
@@
Consider then results for estimates of structural regression coefficients. First, in cases B1, C1, D, E and F (and B2 and B3 when $R^{2}_{\eta}=0$) the estimation model is correctly specified. These results (and those for the cases discussed below) are reported in supplementary Tables A.3–A.25. The broad conclusion is that most of these results are similar to the results in case A. Both estimates perform mostly well, and two-step estimates at least as well as one-step estimates. These conclusions are thus not substantially affected when the models are larger than in case A, in having ordinal items, more latent variables, or non-simple measurement models. This is unsurprising, given that the large-sample properties of the estimators (as discussed in Sections 3.2 and 3.4.1) should not be affected by the complexity of the model.
@@
Even when one-step estimation converges without problems, it is more time-consuming; most of the computing time in these simulations was spent on it. It is likely that this is the most important practical difference between these two methods in large models, where one-step estimation quickly becomes more and more demanding.
@@
The conclusions from these simulations are broadly similar to previous ones for other types of models by Bakk and Kuha (2018) and Rosseel and Loh (2022). In all of these contexts two-step and one-step estimates mostly perform very similarly. Differences between them emerge in difficult situations where the measurement models are weak and sample sizes are small. Here there is some variation between the findings of the different studies. In the latent class examples of Bakk and Kuha (2018), two-step estimation does relatively poorly in difficult settings, possibly because an observed covariate or response provides particularly useful stabilisation for one-step estimates there. In contrast, in the simulations of structural equation models by Rosseel and Loh (2022) it is one-step estimation that performs much less well in the most difficult settings. In our simulations the differences in the hardest situations were somewhat smaller than in these previous studies. It should be noted that the exact simulation settings are not easily comparable. For example, Rosseel and Loh (2022) considered models with even more latent factors (five rather than three), which may be particularly challenging for the small-sample performance of one-step estimation (it would also be computationally very demanding for it for latent trait models). They also considered various cases where the analysis models were misspecified. The conclusions were qualitatively the same as in our simulations, i.e. that both estimates were then biased but the two-step estimates often somewhat less so.
@@ 4.3 Comments on three-step estimates in the simulations
The simulation tables include also results for naive three-step estimates which replace the latent variables $\eta_{ik}$ with their empirical Bayes predictions $\tilde{\eta}_{ik}$. They are typically badly biased for the true structural parameters. This to be expected, based on the general results discussed in Sections 3.4.2 and 3.4.3. These naive estimates are clearly an inadequate way of estimating the structural model for $\eta_{ik}$. As discussed in Section 3.4.2, it is more constructive to regard them as estimates for a model for $\tilde{\eta}_{ik}$, treated as a model of interest in its own right.
An exception here is case C, where the only latent variable $\eta_{1}$ appears as the only explanatory variable in the model. As observed in Section 3.4.3, naive three-step estimation which uses the empirical Bayes predictions then becomes a form of adjusted three-step estimation which avoids measurement error bias. Accordingly, estimates of $\beta_{1}$ in cases C1 and C2 are essentially unbiased and very similar to the two-step and one-step estimates (see supplementary Tables A.11 and A.13). This is not the case for the other simulations here, and generalising this type of adjustment to them would not be as easy.
[…]
@@ 5 Application: Extrinsic and intrinsic work values
In the notation of Section 2, we consider here the structural model $p(\boldsymbol{\eta}_{i(1)}|\mathbf{X}_{i};\boldsymbol{\theta}_{21})$ where $\boldsymbol{\eta}_{i(1)}=(\eta_{i1},\eta_{i2})^{\prime}$ and $\eta_{i1}$ stands for extrinsic and $\eta_{i2}$ for intrinsic work values. They are responses to covariates $\mathbf{X}_{i}$, and there are no observed response variables. As defined in (3), the model includes separate linear models for $\eta_{i1}$ and $\eta_{i2}$, plus their conditional association which we report as the correlation $\rho=\text{corr}(\eta_{i1},\eta_{i2}|\boldsymbol{X}_{i})$. The two latent variables are measured by the corresponding 3 items each, with logistic measurement models (8).
@@
* Model (1) includes only gender as covariate, Model (2) also respondent’s age, partnership status and age of
    the youngest person in the household, and Model (3) also respondent’s education, social class and work status,
    and their parents’ education.
@@
Consider then the estimated structural models for the Netherlands. In terms of the settings of the simulations in Section 4, these data have $n=1374$ and $R^{2}_{\eta}$ around 0.1 for both latent factors, each of them measured by $p=3$ items with $R^{2}_{Y}$ of around 0.3–0.8 and $\pi_{Y}$ of 0.4–0.7 (different for different items). These place this example around the middle of the simulation settings in terms of the strength of the measurement model, but with a somewhat larger sample size than even the larger $n=1000$ considered there. The simulations suggest that the two-step and one-step estimates should both behave well here. They are also similar to each other, with somewhat higher standard errors for the two-step estimates. For the naive three-step estimates, for better comparability we first re-scaled the factor scores so that their variances equal the estimated marginal variances of the corresponding latent variables from step 1 of two-step estimation. Even after this, the three-step estimates of regression coefficients are attenuated toward zero, as would be expected given the theoretical and simulation results above.
@@
Substantively, the results in Table 3 show that individuals’ intrinsic and extrinsic work values, as measured by these survey items, are strongly positively correlated even given covariates. Between genders, there is essentially no difference in extrinsic values, but for intrinsic ones men indicate on average significantly (at the 5% level) higher level of importance on them. Some other significant associations with covariates are also found, as shown by the estimated coefficients in supplementary Tables B.1 and B.2. The importance of extrinsic rewards of a job (pay, hours and holidays) tends to be higher for respondents who are younger, have young children at home, or are currently in paid employment. For intrinsic values (related to the scope of initiative, achievement and responsibility), the expected levels are higher for individuals who have (or whose parents have) higher levels education and who are in professional, administrative or managerial occupations.
@@
Summarising previous literature on gender differences in work values, Bacher et al. (2022) observed that “[…]intrinsic work values, especially altruistic or social work values, seem still to be more important for women than men, at least in some countries and/or in combination with other factors[…]” and that “Gender differences in extrinsic work value orientations are absent or smaller than those for intrinsic work value orientations”. In their own analysis in Austria, they also found higher levels of intrinsic values for women. Our conclusion from the Netherlands is the opposite. This may be because the three intrinsic items in EVS do not include ones of “altruistic or social work” kind. But it could also be an instance of variation between different country contexts. To examine this further, we can use the EVS to carry out a comparative cross-national analysis. Here we include as covariates the respondents’ gender, age, education, social class and work status; the other covariates were omitted from this illustrative example because they had large proportions of missing data in some countries. The list of the 36 countries and their sample sizes in this analysis are shown in Tables B.3 and B.4 of the supplementary appendix.
@@
We allow all parameters of the structural model (regression coefficients and residual variances and covariance) to have different values in each country. Every parameter of the measurement model, in contrast, is constrained to have the same value in all countries, to specify full measurement equivalence across the countries. We note that because the step-1 estimates are here based on much more data than the step-2 estimates (with nearly 59,000 observations in the pooled data but 1000–2000 in each country), step 1 contributes essentially nothing to the standard errors of the estimated structural coefficients. So in this example we could safely treat the measurement parameters from step 1 as known and omit the term for them from the estimated standard errors.
@@
As mentioned at the start of this this section, this application example also allowed us to examine two topics in the relative performance of two-step estimation. From the analysis for the Netherlands above, we discussed the observed level of interpretative confounding in one-step estimates. The second, cross-national analysis provides an extreme illustration of the difference in computational convenience between two- and one-step estimation. The model has 836 estimable parameters. One-step estimates have to be calculated for all of them at once, from the pooled dataset of all the countries. On our computer this took 4.5 hours, using just one starting value. Two-step estimation, in contrast, splits the task in a way which considerably simplifies it. In its step 1, the measurement models are estimated from the pooled dataset, as two one-trait models for three binary items each and without any covariates. Step 2 can then be done separately for each country, because their structural parameters are distinct. Even with multiple starting values (Starts=30 10 in Mplus syntax) this took just 1-2 minutes for both step 1 and each country in step 2, for a total of a little over 0.5 hours for the whole model.

### v3

@@ 3.7 Two-step estimation for other types of models
Two-step estimation could be applied to any instance of the general latent variable model that was defined at the start of Section 2. This requires only that the joint model is itself identified, i.e. that one-step estimates could be calculated. To see this, note first that there are then also identified models for estimating the measurement model in step 1 (in the extreme, the redundant choice of the joint model itself could be used for this). Second, if the structural model is identified when the parameters of the measurement model are estimated with it, it is also identified when these parameters are fixed as in the step 2 of two-step estimation.
@@
To provide a clear focus for this paper, however, we concentrate on the narrower choice of the family of latent trait models that was introduced in Section 2. Before proceeding to empirical investigations of them, we comment briefly on other settings where the two-step method could be used. Some of them are obvious extensions or variants of the models considered here, while others could involve more consequential differences. Implementation and performance of two-step estimation in many of these other cases remain to be examined, but it seems plausible that its behaviour would not be fundamentally different in them.
@@
Considering first the measurement model, some further variants of it would involve no new issues, just appropriate changes to the model specification and the likelihood functions. One obvious class of extensions is produced by models with mixtures of different types of measurement items. In this paper we focus on categorical items, while linear structural equation models (as in Rosseel and Loh 2024) take the items to be continuous. But some applications may involve both of these types, or others such as counts or survival times as measurement items. The latent variables being measured can also be of different types and measured by different types of items, such as when categorical latent variables are measured by categorical items (as in standard latent class analysis) or by continuous ones (the case known as latent profile analysis). Similarly, we could use standard specifications to allow for conditional associations between some measurement indicators, relaxing the assumption that all items in $\mathbf{Y}_{i}$ are conditionally independent given the latent traits $\boldsymbol{\eta}_{i}$.
Other extensions of the measurement model would require further modifications of step 1. As discussed in Section 3.1, the general principle is that this should use the collection of the simplest models that still allow valid estimation of the measurement parameters. We have already noted that one simple instance of this is a situation where two latent variables share part of the measurement model because some items measure both of them (or, similarly, because some indicators of them are conditionally associated). Then step 1 has to be carried out for those two variables together rather than separately. A more involved example of this principle arises for models which include non-equivalence of measurement, i.e. where $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i}^{*};\boldsymbol{\theta}_{1})$ for some subset $\mathbf{X}_{i}^{*}$ of $\mathbf{X}_{i}$. Then the models used for step 1 should also be conditional on $\mathbf{X}_{i}^{*}$ for those elements of $\mathbf{Y}_{i}$ that are affected by the non-equivalence (Vermunt and Magidson 2021a; Lyrvall et al. 2025).
@@
The way to extensions of the structural model is even more obvious. Whatever estimation procedure (and computer code) would be used for one-step estimation of a model is also used, now with measurement parameters fixed, for step 2 of two-step estimation. One simple extension is to models which include mixtures of different types of (continuous and categorical) latent variables. Other generalisations of the structural model could also include such possibilities as interactions or nonlinear terms of the latent variables, or non-recursive models. Two-step estimation in these cases has not been examined. It could behave differently there, if only because such models are inherently more complex and difficult howsoever they are estimated.
@@ 4.1 First set of simulations
We have used simulation studies to examine the properties of two-step estimates for the types of models considered in this paper. In this section we describe them for one model setting in detail. In Section 4.2 we then summarise further simulations in several other settings.
We examine the performance of the point estimates and accuracy of the estimated standard errors. The main comparison of interest is with one-step estimates of the same models. Since they are full maximum likelihood estimates, they serve as a benchmark whose theoretical properties are well understood. In large samples, one-step estimates for correctly specified models will be approximately unbiased and fully efficient, and it is then interesting to see if two-step estimates do roughly as well. In smaller samples, either approach could have an advantage. In this section all the models are correctly specified, but in the next we also include simulations with a misspecified measurement model in order to examine the robustness of the estimates in such cases. Another point of comparison is computational complexity, where two-step estimation should have the advantage. The simulation results also include (naive) three-step estimates, but we defer a discussion of results for them to a separate Section 4.3.
@@
where $\eta_{1i}$ and $\eta_{2i}$ are latent variables and $\epsilon_{i}\sim N(0,\sigma^{2}_{\epsilon})$ is the residual. Here $\eta_{1i}\sim N(0,1)$, $\beta_{0}=0$, and $(\beta_{1},\sigma^{2}_{\epsilon})$ are such that marginally $\eta_{2i}\sim N(0,1)$ in all settings. The value of $\beta_{1}$ is set so that the $R^{2}$ statistic for the response $\eta_{i2}$ is $R^{2}=0,0.2$, or $0.4$. We refer to this as simulation case A. Path diagram for it is shown in Figure A.1 in the supplementary appendix.
The latent variable $\eta_{i1}$ is measured by $p$ binary items $Y_{i1j}$, coded with values 0 and 1, and $\eta_{i2}$ by a different set of $p$ binary items similarly, with $p=4$ or $p=8$. All items are conditionally independent of each other given the latent variables. The measurement model is of the form $\text{logit}[P(Y_{ikj}=1|\eta_{ik})]=\tau_{kj}+\lambda_{kj}\eta_{ik}$ for $k=1,2$ and $j=1,\dots,p$. The parameters $(\tau_{kj},\lambda_{kj})$ are estimated as distinct parameters for different $k,j$, but their true values $\tau_{kj}=\tau$ and $\lambda_{kj}=\lambda$ in the data-generating model are equal for all items in a given simulation. The loading $\lambda$ is set with reference to the linear model for a notional continuous latent variable $Y_{ikj}^{*}$ which implies the logistic model for $Y_{ikj}$. Here $R^{2}$ for $Y_{ikj}^{*}$ given $\eta_{ik}$ is $R^{2}_{Y}=\lambda^{2}\mbox{var}(\eta_{ik})/(\lambda^{2}\mbox{var}(\eta_{ik})+\pi^{2}/3)$ where $\mbox{var}(\eta_{ik})=1$, and $\lambda$ is set so that $R^{2}_{Y}=0.4$ or 0.6. The values of $\tau$ are then set so that the marginal probability $\pi_{Y}=P(Y_{ikj}=1)$ is 0.5 or (approximately) 0.8. In the estimation, we constrain $\tau_{k1}=0$ and $\lambda_{k1}=1$ for the first item $Y_{ik1}$ of each latent variable, leaving $\tau_{kj}$ and $\lambda_{kj}$ for the other p1p-1 items estimable. This affects the implied scales of the latent variables, so that the value of $\beta_{1}$ that the estimators should actually be estimating will depend also on the value of $\lambda$ in the data-generating model. These true values of $\beta_{1}$ are shown in the results tables. In each setting we simulate 1000 datasets with $n=200$ or $n=1000$ independent observations $i$. All combinations of $n$, $R^{2}$, $p$, $\pi_{Y}$ and $R^{2}_{Y}$ are considered, resulting in 48 settings. The two-, one- and naive three-step estimates are obtained using Mplus, combined with R functions for process management as discussed in Section 3.5.
@@ 4.2 Further simulations
  • Ordinal measurement items (case D). Structural model is the same as in case A, but each measurement item $Y_{ikj}$ has 4 rather than 2 categories. Measurement models are again derived from a linear model for a latent continuous $Y_{ijk}^{*}$, here resulting in an ordinal logistic measurement model for each $Y_{ijk}$. Its loading parameter $\lambda$ is set as in case A. The measurement intercepts are such that the marginal probabilities of the four categories of each $Y_{ikj}$ are (0.25,0.25,0.25,0.25) when $R^{2}_{Y}=0.6$ and about (0.20,0.30,0.30,0.20) when $R^{2}_{Y}=0.4$ (both of these correspond to the setting $\pi_{Y}=0.5$ in case A).
@@
Consider then results for estimates of structural regression coefficients. They are reported in supplementary Tables A.3–A.25. First, in cases B1, C1, D, E and F (and B2 and B3 when $R^{2}_{\eta}=0$) the estimation model is correctly specified. The broad conclusion is that most of these results are similar to the results in case A. Both estimates perform mostly well, and two-step estimates at least as well as one-step estimates. These conclusions are thus not substantially affected when the models are larger than in case A, in having ordinal items, more latent variables, or non-simple measurement models. This is unsurprising, given that the large-sample properties of the estimators (as discussed in Sections 3.2 and 3.4.1) should not be affected by the complexity of the model in these (or other) respects.
@@
The conclusions from these simulations are broadly similar to previous ones for other types of models by Bakk and Kuha (2018) and Rosseel and Loh (2024). In all of these contexts two-step and one-step estimates mostly perform very similarly. Differences between them emerge in difficult situations where the measurement models are weak and sample sizes are small. Here there is some variation between the findings of the different studies. In the latent class examples of Bakk and Kuha (2018), two-step estimation does relatively poorly in difficult settings, possibly because an observed covariate or response provides particularly useful stabilisation for one-step estimates there. In contrast, in the simulations of structural equation models by Rosseel and Loh (2024) it is one-step estimation that performs much less well in the most difficult settings. In our simulations the differences in the hardest situations were somewhat smaller than in these previous studies. It should be noted that the exact simulation settings are not easily comparable. For example, Rosseel and Loh (2024) considered models with even more latent factors (five rather than three), which may be particularly challenging for the small-sample performance of one-step estimation (it would also be computationally demanding for it for latent trait models). They also considered various cases where the analysis models were misspecified. The conclusions were qualitatively the same as in our simulations, i.e. that both estimates were then biased but the two-step estimates often somewhat less so.
@@ 4.3 Comments on three-step estimates in the simulations
The simulation tables include also results for naive three-step estimates which replace the latent variables $\eta_{ik}$ with their empirical Bayes predictions $\tilde{\eta}_{ik}$. They are typically badly biased for the true structural parameters. This is to be expected, based on the general results discussed in Section 3.4. These naive estimates are clearly an inadequate way of estimating the structural model for $\eta_{ik}$. As noted in Section 3.4.2, it is more constructive to regard them as estimates for a model for $\tilde{\eta}_{ik}$, treated as a model of interest in its own right.
An exception here is case C, where the only latent variable $\eta_{1}$ appears as the only explanatory variable in the model. As observed in Section 3.4.3, naive three-step estimation which uses the empirical Bayes predictions then becomes a form of adjusted three-step estimation which avoids measurement error bias. Accordingly, estimates of $\beta_{1}$ in cases C1 and C2 are essentially unbiased and very similar to the two-step and one-step estimates (see supplementary Tables A.11 and A.13). This is not the case for the other simulations here, and generalising this type of adjustment to them would not be easy.
The three-step estimates perform relatively well in case G, where the estimation model is misspecified because of omitted cross-loadings and where two-step and one-step estimates are clearly biased. The naive three-step estimates avoid this misspecification bias. This is because they only require valid estimates of the measurement models for each latent variable individually, and those are here correctly estimated (even though it includes cross-loadings, the model in case G implies also a one-trait model for $Y_{21}$$Y_{24}$ given $\eta_{2}$). However, the three-step estimates still have their usual measurement error bias in this case.
@@ 5 Application: Extrinsic and intrinsic work values
In the notation of Section 2, the structural model considered here is $p(\boldsymbol{\eta}_{i(1)}|\mathbf{X}_{i};\boldsymbol{\theta}_{21})$, where $\boldsymbol{\eta}_{i(1)}=(\eta_{i1},\eta_{i2})^{\prime}$ and $\eta_{i1}$ stands for extrinsic and $\eta_{i2}$ for intrinsic work values. They are responses to covariates $\mathbf{X}_{i}$, and there are no observed response variables. As defined in (3), the model includes separate linear models for $\eta_{i1}$ and $\eta_{i2}$, plus their conditional association which we report as the correlation $\rho=\text{corr}(\eta_{i1},\eta_{i2}|\boldsymbol{X}_{i})$. The two latent variables are measured by the corresponding 3 items each, with logistic measurement models (8).
@@
* Model (1) includes gender as covariate, Model (2) also respondent’s age, partnership status and age of youngest person
in the household, and Model (3) also respondent’s education, social class and work status, and their parents’ education.
@@
Consider then the estimated structural models for the Netherlands. In terms of the settings of the simulations in Section 4, these data have $n=1374$ and $R^{2}_{\eta}$ around 0.1 for both latent factors, each of them measured by $p=3$ items with $R^{2}_{Y}$ of around 0.3–0.8 and $\pi_{Y}$ of 0.4–0.7 (different for different items). These place this example around the middle of the simulation settings in terms of the strength of the measurement model, but with a somewhat larger sample size than even the larger $n=1000$ considered there. The simulations suggest that the two-step and one-step estimates should both behave well. Here they are also similar to each other, with somewhat higher standard errors for the two-step estimates. For the naive three-step estimates, for better comparability we first re-scaled the factor scores so that their variances equal the estimated marginal variances of the corresponding latent variables from step 1 of two-step estimation. Even after this, the three-step estimates of regression coefficients are attenuated toward zero, as would be expected given the theoretical and simulation results above.
@@
The results in Table 3 indicate that individuals’ intrinsic and extrinsic work values are strongly positively correlated even given covariates. Between genders, there is essentially no difference in extrinsic values, but for intrinsic ones men indicate on average significantly (at the 5% level) higher level of importance on them. Some other significant associations with covariates are also found, as shown by the estimated coefficients in supplementary Tables B.1 and B.2. The importance of extrinsic rewards of a job (pay, hours and holidays) tends to be higher for respondents who are younger, have young children at home, or are currently in paid employment. For intrinsic values (related to the scope of initiative, achievement and responsibility), the expected levels are higher for individuals who have (or whose parents have) higher levels education and who are in professional, administrative or managerial occupations.
@@
Summarising previous literature on gender differences in work values, Bacher et al. (2022) observed that “[…]intrinsic work values, especially altruistic or social work values, seem still to be more important for women than men, at least in some countries and/or in combination with other factors[…]” and that “Gender differences in extrinsic work value orientations are absent or smaller than those for intrinsic work value orientations”. In their analysis in Austria, they also found higher levels of intrinsic values for women. Our conclusion from the Netherlands is the opposite. This may be because the three intrinsic items in EVS do not include ones of “altruistic or social work” kind. But it could also reflect variation between different country contexts. To examine this further, we can use the EVS to carry out a comparative cross-national analysis. Here we include as covariates the respondents’ gender, age, education, social class and work status; the other covariates were omitted from this illustrative example because they had large proportions of missing data in some countries. The list of the 36 countries and their sample sizes in this analysis are shown in supplementary Tables B.3 and B.4.
@@
We allow all parameters of the structural model (regression coefficients and residual variances and covariance) to have different values in each country. Every measurement parameter, in contrast, is constrained to have the same value in all of them, to specify measurement equivalence across the countries. We note that because the step-1 estimates are thus based on much more data than the step-2 estimates (with ca. 59,000 observations in the pooled data but 1000–2000 in each country), step 1 contributes essentially nothing to the standard errors of the estimated structural coefficients. So in this example we could safely treat the measurement parameters from step 1 as known and omit the term for them from the estimated standard errors.
@@
As mentioned at the start of this this section, this application example also allowed us to examine two topics in the relative performance of two-step estimation. For the analysis for the Netherlands above, we discussed the level of interpretative confounding in one-step estimates. The second, cross-national analysis provides an extreme illustration of the difference in computational convenience between two- and one-step estimation. The model has 836 estimable parameters. One-step estimates have to be calculated for all of them at once, from the pooled dataset of all the countries. On our computer this took 4.5 hours, using just one starting value. Two-step estimation, in contrast, splits the task in a way which considerably simplifies it. In its step 1, the measurement models are estimated from the pooled dataset, as two one-trait models for three binary items each and without any covariates. Step 2 can then be done separately for each country, because their structural parameters are distinct. Even with multiple starting values (Starts=30 10 in Mplus syntax) this took just 1-2 minutes for both step 1 and each country in step 2, for a total of a little over 0.5 hours for the whole model.